OpenStudy (anonymous):

Please help: find the real number solutions of the equation: x^3-1331=0 please help!!

OpenStudy (anonymous):

11

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

x^3 -1331 = 0 x^3 = 1331 is now so do cubic root of 3! x = 11 :)

OpenStudy (anonymous):

Thats x, what's the real number solutions

OpenStudy (anonymous):

S = [ x ∈ R / x = 11] that? :s

OpenStudy (anonymous):

?